OpenStudy (anonymous):

Add this to that, divide by three; the square of this of course, you’ll see. But that to this is eight to one; so find what this is and you’re done.

OpenStudy (turingtest):

Let that=x and this=y Add this to that, divide by three; the square of this of course, you’ll see.\[\frac{x+y}3=y^2\]But that to this is eight to one\[x=8y\]so find what this is and you're done\[\frac{8y+y}3=y^2\]